OpenStudy (anonymous):

y=2x^2-7x+12 complete the square to find the vertex

OpenStudy (anonymous):

vertex is at \[-\frac{b}{2a}\] in this case \[\frac{7}{4}\]

OpenStudy (anonymous):

to find the second coordinate replace x by \[\frac{7}{4}\]and get \[\frac{47}{8}\]